the coach wants to spend three-eighths of basketball practice time on defense. He plans to spend 45 min. on this. how many hours will the team spend practicing?
Answer by Theo(13342)

let x equal the amount of time spent practicing.

coach wants to spend 3/8 of that on defense.

he plans to spend 45 minutes on defense.

3/8 * x = 45 minutes.

solve for x to get x = 8/3 * 45 = 8 * 15 = 120 minutes total.

that would be 2 hours.

2 hours is your solution.